(25 points)

he problem below will have you perform some basic calculations related to cost concepts. Space for each calculation is provided below, along with highlighted areas for you to enter your solutions.

McKnight Enterprises sells only a single product. Data from the company’s most recent income statement is provided below. Using that data, complete the calculations required.

Sales (4,000 units)

$120,000

Less variable expenses

(68,000)

Contribution margin

52,000

Less fixed expenses

(40,000)

Net income

$ 12,000

Please enter your answer (no calculations necessary) in the “Answer/Solution” column.

Requirement

Answer/Solution

Points

Possible

The contribution margin per unit is:

$

5

If sales doubled to $240,000, then the total variable costs would equal:

$

5

If sales doubled to $240,000, then the total fixed costs would equal:

$

5

Compute the number of units that must be sold to break even.

5

Compute the number of units that must be sold to achieve a profit of $20,000.

5
Answer #1

Question - 1

Contribution Margin per Unit = Selling price per Unit - Variable cost per Unit = 120,000 / 4000 - 68000 / 4000

= 30 - 17 = 13 per Unit

Question - 2

When sales double, variable cost also doubles = 2 * 68000 = 136000

Question - 3

Fixed cost remains constant, even if sales doubles. Hence Fixed cost equal ............40000

Question - 4

Number of units needed to break even = Fixed cost / Contribution margin per Unit = 40000 / 13 = 3077 Units

Question - 5

Number of units needed to achieve a profit of 20000

= (Fixed cost + Required profit) / Contribution margin per Unit = (40000 + 20000) / 13 = 4616 Units